1.calculate the net ATPs produced from the pyruvate through the TCA cycle.


2.calculate the net ATPs produced from one molecule of glucose in a glycolytic pathway

1
Expert's answer
2022-03-28T17:06:02-0400

1. 

Stage1: ATP=2ATP

2NADH+H+-------->2FADH2(to ETC)=3ATP

Stage2: conversion of yruvate to ACoA

2NADH +H+(to ETC) =5ATP

Stage3: TCA cycle

ATP(at one site)=2ATP

NADH+H+at three syeps(to ETC)=15ATP

FADH2 at one step(to ETC)=3ATP

Total ATP from one molecule of glucose=2+3+5+2+15+3=30



2. Since in whole processthe

: Total ATP consumed= 2


: Total ATP formation=2×2=4 formation of NADH+H=2


: And we know that 1 NADH2= 3ATP


: Therefore net ATP synthesis=


: 4+2×3–2=8 molecules of ATP